OpenStudy (amistre64):

converge or diverge? and how to prove it: \[\sum_{1}^{inf} (-1)^{n-1}\frac{2^n}{n^4}\]

OpenStudy (amistre64):

I had tried, but not confident enough to know if im right :) \[\lim_{n->inf}\ \sqrt[n]{\left|\frac{2^n}{n^4}\right|}\] \[\lim_{n->inf}\ \frac{2}{n^{4/n}}\] \[\frac{2}{n^{0}}=2\] \[2\cancel\lt1\] diverges

OpenStudy (mr.math):

What do you think about \(\large \lim_{n \to \infty} \frac{2^n}{n^4}?\)

OpenStudy (amistre64):

isnt that inf/inf? are we good to lhop it?

OpenStudy (mr.math):

I think so!

OpenStudy (amistre64):

2^n (ln(2))^4 ----------- might be the lhop; whic is inf 4!

OpenStudy (mr.math):

If you want to apply L'Hopital"s rule, you will have to do it 4 times in order to get rid of the n in the denominator. For me, it's clear that 2^n grows much faster than n^4 and therefore the limit is infinity.

OpenStudy (amistre64):

yeah, the blurb at the start got me on that one; but then the limit doesnt matter about the start, it only cares about the farout

OpenStudy (amistre64):

mighta been a different blurb on another problem im thinking of :)

OpenStudy (mr.math):

I don't get what you mean. Your English is a little bit difficult for me, sorry! :D

OpenStudy (amistre64):

thats ok, i resolved my own quandry :) thnx

OpenStudy (mr.math):

You're welcome :)

OpenStudy (mr.math):

Oh and being not able to understand you is my problem, I need to work on my English. Just had to say that!!

OpenStudy (anonymous):

\[2^x\] grows faster than any polynomial in x

OpenStudy (zarkon):

you really cant write this \[\frac{2}{n^{0}}=2\] if you took the limit there is not n

OpenStudy (zarkon):

*should be no n

OpenStudy (anonymous):

\[\lim_{n\rightarrow \infty}n^{\frac{4}{n}}=\lim_{n\rightarrow \infty}e^{\frac{4log(n)}{n}}=1\]should work